X+2 over 3= -1/3[X+2]
Answer by checkley71(8403)

(X+2)/3=-1/3(X+2)
(X+2)/3=-(X+2)/3 CROSS MULTIPLY
3(X+2)=3*-(X+2)
3X+6=-3X-6
3X+3X=-6-6
6X=-12
X=-12/6
X=-2 ANSWER.
PROOF:
(-2+2)/3=-1/3(-2+2)
0/3=-1(0)/3
0/3=0/3